What will be the integration of dx/(3sinx + 4cosx)?

Grade:12

1 Answers

Astro Reporter
7 years ago
Int dx/(3sinx + 4cosx) = Int dx/[5(3/5sinx + 4/5cosx)]Since, sin37=3/5 and cos37=4/5, therefore the above eq becomes,=Int dx/5(sin37sinx + cos37cosx)=Int dx/5cos(x-37)=Int 1/5•sec(x-37)dx=1/5•log|sec(x-37)+tan(x-37)|+c